OpenStudy (anonymous):

Express the complex number in trigonometric form. 3 3(cos 0° + i sin 0°) 3(cos 90° + i sin 90°) 3(cos 180° + i sin 180°) 3(cos 270° + i sin 270°)

OpenStudy (anonymous):

I'm not quite sure I understand your question, because those numbers already are in trigonometric form. The only thing I can say, is that it can be abbreviated as 3cis(0°), 3cis(90°), 3cis(180°), 3cis(270°)

OpenStudy (anonymous):

@kikazo those are the answer choices under the 3.

OpenStudy (anonymous):

So your question is how to write it as cis? Well, just look at the angle it has. For instance, 3(cos(0°)+i sin(0°))=3cis(0°), and so on. If you look carefully, the angle is the same in both sin and cos for each of them, so you just have to change the whole parenthesis by cis(angle). The coefficient stays the same.

OpenStudy (anonymous):

@kikazo so which one is the answer?

OpenStudy (anonymous):

\[ \cos 0^\circ = 1, \sin 0^\circ= 0 \]

OpenStudy (anonymous):

3(cos 0° + i sin 0°) = 3(1) = 3

OpenStudy (anonymous):

Oh, wait, I just understood the question! xD sorry, yea, wio is right